# CodeGRAG: Bridging the Gap between Natural Language and Programming Language via Graphical Retrieval Augmented Generation

### v3

@@ Abstract
Utilizing large language models to generate codes has shown promising meaning in software development revolution. Despite the intelligence shown by the general large language models, their specificity in code generation can still be improved due to the syntactic gap and mismatched vocabulary existing among natural language and different programming languages. In this paper, we propose CodeGRAG, a Graphical Retrieval Augmented Code Generation framework to enhance the performance of LLMs. CodeGRAG builds the graphical view of code blocks based on the control flow and data flow of them to fill the gap between programming languages and natural language, which can facilitate natural language based LLMs for better understanding of code syntax and serve as a bridge among different programming languages. To take the extracted structural knowledge into the foundation models, we propose 1) a hard meta-graph prompt template to transform the challenging graphical representation into informative knowledge for tuning-free models and 2) a soft prompting technique that injects the domain knowledge of programming languages into the model parameters via finetuning the models with the help of a pretrained GNN expert model. CodeGRAG significantly improves the code generation ability of LLMs and can even offer performance gain for cross-lingual code generation. Implementation is available at https://anonymous.4open.science/r/Code-5970/ .
CodeGRAG: Bridging the Gap between Natural Language and Programming Language via Graphical Retrieval Augmented Generation
Kounianhua Du1, Jizheng Chen1, Renting Rui1, Huacan Chai1, Lingyue Fu1, Wei Xia2, Yasheng Wang2, Ruiming Tang2, Yong Yu1, Weinan Zhang1 1Shanghai Jiao Tong University, 2 Huawei Noah’s Ark Lab Shanghai, China {kounianhuadu, wnzhang}@sjtu.edu.cn
@@ 1 Introduction
In recent years, large language models (LLMs) (Achiam et al., 2023; Touvron et al., 2023a) have shown great impact in various domains. Automated code generation emerges as a captivating frontier (Zheng et al., 2023; Roziere et al., 2023; Shen et al., 2023), promising to revolutionize software development by enabling machines to write and optimize code with minimal human intervention.
However, syntactic gap and mismatched vocabulary among natural language (NL) and programming languages (PL) exist, hindering LLM’s performance on code generation. As illustrated in Figure 1, programming language (marked in blue) contains special tokens such as “int” or “++” that natural language (marked in yellow) doesn’t possess, leading to vocabulary mismatch. Besides, the relations between tokens in programming languages are often structural, e.g., the complex branching and jumps, whereas natural language is arranged simply in sequential manner, leading to syntactic gap. For example, in the control flow graph of the raw code (marked in pink), two “if” blocks (marked in purple) are adjacent and are executed sequentially under certain condition, but they appear to be intervaled in raw textual code.
As discussed above, the innate structures of programming languages are different from that of the sequential-based natural language. The challenges of enhancing a general-purposed large language models for code-related tasks can be summarized into two folds.
@@
(C1) How to solve the gap between different languages and better interpret the inherent logic of code blocks. Code, unlike natural language, possesses a well-defined structure that governs its syntax and semantics. This structure provides valuable information about the relationships between different parts of the code, the flow of execution, and the overall organization of the functions Jiang et al. (2021); Guo et al. (2020). General-purpose LLMs regard a code block as a sequence of tokens. By ignoring the inherent structure of codes, they miss out on essential cues that could help them better understand and generate code. In addition, the multi-lingual code generation abilities of LLMs is challenging due to the gap among different programming languages.
(C2) How to inject the innate knowledge of programming languages into general purpose large language models for enhancement. Despite the well representation of the programming knowledge, the ways to inject the knowledge into the NL-based foundation models is also challenging. The structural representation of codes could be hard to understand, which poses a challenge to the capability of the foundation models.
To solve the above challenges, we propose CodeGRAG, a graphical retrieval augmented generation framework for code generation. For (C1), we propose to interpret the code blocks using the composed graph based on the data-flow and control-flow of the code block, which extracts both the semantic level and the logical level information of the code. The composed graphical view could 1) better capture the innate structural knowledge of codes for NL-based language models to understand and 2) model the innate function of code blocks that bridging different programming languages. For (C2), we propose a meta-graph prompting technique for tuning-free models and a soft-prompting technique for tuned models. The meta-graph prompt summarizes the overall information of the extracted graphical view and transforms the challenging and noisy graphical representation into informative knowledge. The soft-prompting technique deals with the graphical view of codes with a pretrained GNN expert network and inject the processed knowledge embedding into the parameters of the general-purpose foundation models with the help of supervised finetuning.
@@
  • Novel GraphRAG framework for code generation. We propose CodeGRAG that bridges the gap among natural language and programming languages, transfers knowledge among different programming languages, and enhances the ability of LLMs for code generation. CodeGRAG requires only one calling of LLMs and can offer multi-lingual enhancement.
  • Effective graphical view to inform and stimulate the structural programming knowledge of LLMs. We propose an effective graphical view to purify the semantic and logic knowledge from the code space, which offers more useful information than the raw code block and can summarize the cross-lingual knowledge.
  • Effective soft prompting technique to preserve the programming domain knowledge and inject it into LLMs parameters. We propose an effective soft prompting technique, which injects the domain knowledge of programming languages into the model parameters via finetuning LLMs with the assistance of a pretrained GNN expert model.
@@ 2 Methodology
Figure 2: Overview of CodeGRAG. (Top) Knowledge Preparation. We extract the control flow and data flow of each external code block and compose them using the read-write signal to obtain the semantic and logical expression of each code block, which is then abstracted into graphical view as hard knowledge document and embedded into GraphEmb as soft knowledge document. The GraphEmb is encoded by a pretrained GNN expert model constrained by the alignment and structure preserving objectives. (Bottem) Retrieval Augmented Generation. We extract query from the task input and retrieve from the external corpus. For tuning free models, we use the hard graphical view to stimulate the structural programming knowledge of LLMs for enhanced generation. For tunable models, we use the soft GraphEmb and inject the programming domain knowledge into LLMs parameters via finetuning them with the GNN expert signals. The expert signals informed LLMs can then produce enhanced generation.
@@ 2.1 Overview
In this paper, we leverage both generative models and retrieval models to produce results that are both coherent and informed by the expert graphical knowledge of programming language. The overall process of CodeGRAG is illustrated in Figure 2, which mainly consists of three stages: graphical knowledge base preparation, knowledge querying, and graphical knowledge augmented generation.
@@ 2.2 Graphical Knowledge Base Preparation
In this section, we discuss how to extract informative graphical views for code blocks. We analyze the syntax and control information of code blocks and extract their graphical views to better represent the codes. This process can be formulated as, $\forall c_{i}\in D_{\text{pool}}$:
$g_{i}\xleftarrow{}\text{GraphExtractor}(c_{i}),$ (1)
[…]
where $c_{i}$ is the raw code block and $g_{i}$ is the corresponding extracted graphical view.
To capture both the semantic and the logical information, we propose to combine the data flow graph (Aho et al., 2006) and the control flow graph (Allen, 1970) with the read-write signals (Long et al., 2022) to represent the code blocks, both of them are constructed on the base of the abstract syntax tree.
@@
An illustration of the composed graphical view is in Figure 3. After obtaining the composed syntax graphs, we use them to inform the general-purpose LLMs to bridge the gap between NL and PLs, where both the semantic level and the logic level information are preserved.
2.3 Knowledge Querying
Given a target problem to be completed, we generate informative query of it and use it to retrieve graphical knowledge from the constructed knowledge base.
We extract the problem description of each task to reduce the ambiguity and then concatenate it with the function declaration to serve as the query content, where the functionality and input format of the expected code block are contained. The query of the retrieval includes problem description $Q_{p}$ and function description $Q_{c}$, while each content of the retrieval pool includes raw code block $V_{c}$ and its graphical view $V_{g}$.
To expressively represent the components, we use the encoder $\phi(\cdot)$ of the pretrained NL2Code model to represent the problem description and code snippets. The retrieval function is:
$\mathbf{h^{V}}$ $=\phi(V_{c}\|V_{g}),$ (3)
$\mathbf{h^{Q}}$ $=\phi(Q_{p}\|Q_{c}),$ (4)
Distance $=1-\frac{\mathbf{\mathbf{h^{Q}}\cdot\mathbf{h^{V}}}}{\|\mathbf{h^{Q}}\|\cdot\|\mathbf{h^{V}}\|}.$ (5)
2.4 Graphical Knowledge Augmented Generation
After we obtain the returned graphical view, we inject it to the foundation LLMs for graphical knowledge augmented generation. Since the graphical view is hard to understand, we propose 1) a meta-graph template to transform the graphical view into informative knowledge for tuning-free model and 2) a soft prompting technique to tune the foundation models for their better understanding of the graphical views with the assistance of an expert GNN model.
@@ 2.4.1 Hard Meta-Graph Prompt
The original graphical view of a code block could contain hundreds of nodes and edges. A full description of it could cost a overly long context, along with the understanding challenge posed by the long edge lists. Therefore, we propose to use a meta-graph template to abstract the information of the graphical view. The abstracted meta-graph consists of the canonical edge types and node types, which describes the basic topology of the graphical view (Sun and Han, 2013), with the textual features obtained from the ASTs contained in the node and edge features.
Then we use the meta-graph template to transform the retrieved graphical view into digestable knowledge and insert it into the final prompt for generation. As illustrated in Figure 5 in the Appendix, the final prompt consists of three components: the system prompt illustrated in the blue part, the retrieved knowledge and hints illustrated in the green part, and the problem (including task description, function declaration, etc.) illustrated in the yellow part. The three parts are concatenated to be fed into LLMs for knowledge augmented generation.
@@ 2.4.2 Soft Prompting with the Expert
Directly hard prompt to the LLMs poses a challenge to the digesting capability of the backbone LLMs, which could fail under the case where the backbone LLMs cannot well understand the graph components. To compress the graphical knowledge into model parameters and help the backbone LLMs to better understand the programming language, we propose a soft prompting technique. The overall procedure can summarized into expert encoding of graphical views, finetuning with the expert signal, and inference.
Expert Encoding of Graphical Views. We design a graph neural network to preserve the semantic and logical information of code blocks. The representation of each node $\mathbf{n}_{i}^{(0)}$ and edge $\mathbf{e}_{i}^{(0)}$ are first initialized with vectors corresponding to the node text and edge text encoded by $\phi_{1}$. A message passing process is first conducted to fuse the semantic and structural information into each node representation.
@@
The expert encoding network is optimized via the contrastive learning based self-supervised training, which includes the intra-modality contrastive learning and inter-modality contrastive learning. The intra-modality constrastive learning serves for preserving the modality information, while the inter-modality contrastive learning serves for modality alignment.
  • Alignment Contrastive Learning. There are two types of alignment to be ensured: 1) NL-Code (NC) alignment and 2) Code-Graph (CG) alignment. We define the positive pairs for NC alignment purpose as $\mathcal{I}^{+}_{NC}=\{\langle\mathbf{h}_{i}^{V},\mathbf{h}_{i}^{Q}\rangle|i\in D_{\text{train}}\}$ and define the negative pairs for NC alignment purpose as $\mathcal{I}^{-}_{NC}=\{\langle\mathbf{h}_{i}^{V},\mathbf{h}_{j}^{Q}\rangle|i\neq j,i\in D_{\text{train}},j\in D_{\text{train}}\}.$
    And we define the positive pairs for CG alignment purpose as $\mathcal{I}^{+}_{CG}=\{\langle\phi_{1}(c_{i}),\phi_{2}(g_{i})\rangle|i\in D_{\text{train}}\}$ and define the negative pairs for CG alignment purpose as $\mathcal{I}^{-}_{CG}=\{\langle\phi_{1}(c_{i}),\phi_{2}(g_{j})\rangle|i\neq j,i\in D_{\text{train}},j\in D_{\text{train}}\}.$
  • Structure Preserving Contrastive Learning. To preserve the structural information of the graphical views, we perform intra-modality contrastive learning among the graphical views and their corrupted views. Concretely, we corrupt each of the graphical view $g_{i}$ with the edge dropping operation to obtain its corrupted view $g_{i}^{\prime}$. The positive pairs for structure-preserving purpose are then designed as $\mathcal{I}^{+}_{\text{preserve}}=\{\langle\phi_{2}(g_{i}),\phi_{2}(g_{i}^{\prime})\rangle|i\in D_{\text{train}}\}$. The negative pairs for structure preserving purpose are designed as $\mathcal{I}^{-}_{\text{preserve}}=\{\langle\phi_{2}(g_{i}),\phi_{2}(g_{j}^{\prime})\rangle|i\neq j,i\in D_{\text{train}},j\in D_{\text{train}}\}$.
Finetuning with the Expert Soft Signal. To help the backbone LLMs to digest the graphical views, we tune the LLMs with the expert soft signal using supervised finetuning. The prompt for finetuning consists of the system prompt, retrieved knowledge where the expert encoded graphical view is contained using a token embedding, and task prompt, which is illustrated in Figure 5 in the Appendix.
Inference. After the finetuning stage, we used the tuned models to generate codes using the soft prompting template as illustrated in Figure 5 in the Appendix.
@@ 3 Experiments
    Can soft prompting enhance the capability of the backbone LLMs? Does finetuning with the soft prompting outperforms the simple supervised finetuning?
@@
    How is the compatibility of the graphical view?

### v4

@@ Abstract
Utilizing large language models to generate codes has shown promising meaning in software development revolution. Despite the intelligence shown by the large language models, their specificity in code generation can still be improved due to the syntactic gap and mismatched vocabulary existing between natural language and programming languages. In this paper, we propose CodeGRAG, a Graphical Retrieval Augmented Code Generation framework that bridges the gap between NL and PL to enhance the performance of LLMs. CodeGRAG builds the graphical view of code blocks based on the control flow and data flow of them to better interpret the programming domain knowledge, which can facilitate natural language based LLMs for better understanding of code syntax and serve as a bridge among different programming languages. To take the extracted structural knowledge into the foundation models, we propose 1) a hard meta-graph prompt template to transform the challenging syntax graph into informative graphical view for tuning-free models and 2) a soft prompting technique that injects the domain knowledge of programming languages into model parameters via finetuning the models with the soft signals encoded by GNN expert model. Specifically, two constraints are designed to improve the alignment and structure expressiveness, contributing to the informativeness of the single-token-sized external <GraphEmb> for enhanced code generation. CodeGRAG significantly improves the code generation ability of LLMs and can even offer performance gain for cross-lingual code generation. Implementation is available at https://anonymous.4open.science/r/Code-5970/ .
@@ 1 Introduction
In recent years, large language models (LLMs) (Achiam et al., 2023; Touvron et al., 2023) have shown great impact in various domains. Automated code generation emerges as a captivating frontier (Zheng et al., 2023; Roziere et al., 2023; Shen et al., 2023), promising to revolutionize software development by enabling machines to write and optimize code with minimal human intervention.
Code generation involves translating natural language instructions into structured and executable code in a target programming language. However, syntactic gap and mismatched vocabulary between natural language (NL) and programming languages (PL) exist, hindering LLM’s performance on code generation. As illustrated in Figure 1, programming language (marked in blue) contains special tokens such as “int” or “++” that natural language (marked in yellow) doesn’t possess with the same physical significance, leading to vocabulary mismatch. Besides, the relations among tokens in programming languages are often structural, e.g., the complex branching and jumps, whereas natural language is arranged simply in sequential manner, leading to syntactic gap. For example, in the control flow graph of the raw code (marked in pink), two “if” blocks (marked in purple) are adjacent and are executed sequentially under certain condition, while appearing distant in the raw textual code.
As discussed above, the challenges of enhancing large language models for code-related tasks can be summarized into two folds.
(C1) Effective representation of programming hints to interpret the inherent logic. Code, unlike natural language, possesses a well-defined structure that governs its syntax and semantics. This structure provides valuable information about the relationships between different parts of the code, the flow of execution, and the overall organization of the functions (Jiang et al., 2021; Guo et al., 2020). LLMs regard a code block as a sequence of tokens. By ignoring the inherent structure of codes, they miss out on essential cues that could help them better understand and generate code. Previous RAG methods (Lu et al., 2022; Su et al., 2024; Zhou et al., 2022) stimulate the code generation abilities of LLMs by using relevant code snippets or document as hints, failing to well represent the innate structures.
@@
(C2) Effective ways to digest the programming domain knowledge. In addition to the well representation of the programming knowledge, the ways to inject the knowledge into the NL-based LLMs is also challenging. The structural representation of codes and the unfamiliar vocabulary could be hard to understand, which poses a challenge to the capability of the foundation models.
To solve the above challenges, we propose CodeGRAG, a graphical retrieval augmented generation framework for code generation. For (C1), we propose to interpret the code blocks using the graphical view composed from data-flow and control-flow of the code block to serve as the programming hints, which extracts both the semantic level and the logical level information of the code. The composed graphical view could 1) better capture the innate structural knowledge of codes for NL-based language models to understand and 2) model the innate function of code blocks that bridge different programming languages. For (C2), we propose a hard meta-graph prompting technique for tuning-free models and a soft-prompting technique for tunable models. The hard meta-graph prompt summarizes the overall information of the extracted graphical view and transforms the challenging and noisy graphical representation into informative knowledge. The soft-prompting technique injects the programming domain knowledge into LLMs by tuning them with soft signals from expert GNN-encoded graphical views, under the regularization of alignment and structure-preserving objectives. After the tuning, LLMs are capable to well digest the structural graphical views by taking in the soft <GraphEmb> at single-token cost, avoiding them from the vocabulary mismatch and syntactic gap problems.
@@
  • Identification of the vocabulary mismatch and syntactic gap problems in NL2Code process. We explain the identified problems with explicit examples and propose CodeGRAG, a novel graphical retrieval augmented generation framework, to tackle them for enhanced code generation.
  • Effective graphical view to inform and stimulate the programming knowledge of LLMs. We propose an effective graphical view to serve as the programming hints, which offers more informative knowledge than the raw code block.
  • Effective soft prompting technique to inject the programming domain knowledge into LLMs and help them better digest the programming hints. We propose a soft prompting technique, which injects the domain knowledge of programming languages into the model parameters via finetuning LLMs with the soft signals of GNN-encoded graphical representations. A two-hop contrastive objective for alignment along with a graph contrastive objective for structure preserving are designed to improve the expressiveness of the soft signal.
@@ 2 Methodology
Figure 2: Overview of CodeGRAG. (Top) Knowledge Preparation. We extract composed syntax graphs of external code blocks by composing the control flow and data flow of codes using the read-write signal, preserving the innate semantic and logical information. The composed graphs are then abstracted into graphical views as hard knowledge document and embedded into <GraphEmb>s as soft knowledge document. The <GraphEmb> is encoded by a pretrained GNN expert model constrained by the alignment and structure preserving objectives. (Bottom) Retrieval Augmented Generation. We extract query from the task input and retrieve from the external corpus. For tuning free models, we use the hard graphical view to stimulate the structural programming knowledge of LLMs for enhanced generation. For tunable models, we use the soft <GraphEmb> and inject the programming domain knowledge into LLMs parameters via finetuning them with the GNN expert signals. The expert signals informed LLMs can then produce enhanced generation.
@@ 2.1 Overview
In this paper, we leverage both generative models and retrieval models to produce results that are coherent and informed by expert graphical knowledge of the programming language. The overall process of CodeGRAG is illustrated in Figure 2.
@@ 2.2 Graphical Knowledge Base Preparation
To capture both the semantic and the logical information, we represent code blocks by combining the data flow graph (Aho et al., 2006) and the control flow graph (Allen, 1970) with the read-write signals (Long et al., 2022), both of them are constructed on the basis of the abstract syntax tree.
@@
An illustration of the composed graphical view is in Figure 3. After obtaining the composed syntax graphs, we use them to inform the LLMs to bridge the gap between NL and PLs, where both the semantic level and the logic level information are preserved.
2.3 Graphical Retrieval Augmented Generation
In this section, we discuss the ways to effectively utilize the constructed knowledge base.
For the retrieval algorithm, we use CodeT5 to encode the query and corpus, measuring cosine distances among vectors to obtain the top-1 relevant document as the knowledge content. Since the focus of our paper is to discuss the representation of the used knowledge content but not the accuracy of ranking, we keep ranking algorithm the same for all used methods.
For the representation of the retrieved knowledge content, since the composed syntax graph is hard to understand, we propose 1) a hard meta-graph template to transform the challenging knowledge into informative graphical view for tuning-free model and 2) a soft prompting technique to tune the foundation models for their better understanding of the graphical views with the assistance of an expert GNN model.
@@ 2.3.1 Hard Meta-Graph Prompt
The original composed syntax graph of a code block could contain hundreds of nodes and edges. A full description of it could cost an overly long context, along with the understanding challenge posed by the long edge lists. Therefore, we propose to use a meta-graph [^1] https://www.dgl.ai/dgl_docs/generated/dgl.DGLGraph.metagraph.html to abstract the original graph into informative graphical view. The abstracted meta-graph consists of the canonical edge types and node types, which describes the basic topology of the composed syntax graph (Sun and Han, 2013), with the textual features obtained from the ASTs contained in the node and edge features.
The template for the meta-graph is displayed as below. {mdframed} Graph(num_nodes={$node\_type:\#nodes$},
    num_edges=
{$(src\_node\_type,edge\_type,$ $dst\_node\_type):\#edges$},
[…]
Then we use the meta-graph template to transform the retrieved graphical view into digestable knowledge and insert it into the final prompt for generation. As illustrated in Figure 4 (a), the final prompt consists of three components: the system prompt illustrated in the blue part, the retrieved knowledge and hints illustrated in the green part, and the problem (including task description, function declaration, etc.) illustrated in the yellow part. The three parts are concatenated to be fed into LLMs for knowledge augmented generation.
Figure 4: Prompt templates.
@@ 2.3.2 Soft Prompting with the Expert
Directly hard prompt to the LLMs poses a challenge to the digesting capability of the backbone LLMs, suffering from the vocabulary mismatch and syntactic gap problems. To compress the graphical knowledge into model parameters and help the backbone LLMs to better understand the programming language, we propose a soft prompting technique. The overall procedure can be summarized into expert encoding of graphical views, finetuning with the expert signal, and inference.
We design a graph neural network to preserve the semantic and logical information of code blocks. The representation of each node $\mathbf{n}_{i}^{(0)}$ and edge $\mathbf{e}_{i}^{(0)}$ are first initialized with vectors corresponding to the node text and edge text encoded by $\phi_{1}$, which is a specialized coder model trained to align programming and natural languages. A message passing process is first conducted to fuse the semantic and structural information into each node representation.
@@
The expert encoding network is optimized via the contrastive learning based self-supervised training. A two-hop contrastive learning objective is designed to align the code-NL-PL spaces, while a structure preserving contrastive learning is conducted to improve the expressiveness of the expert GNN and avoid the aligning process from being dominated by the NL side.
  • Two-Hop Alignment Contrastive Learning. There are two types of alignment to be ensured: 1) NL-Code (NC) alignment and 2) Code-Graph (CG) alignment.
    We define the positive pairs and negative pairs for NC alignment purpose as:
    $\mathcal{I}^{+}_{NC}=\{\langle\mathbf{h}_{i}^{V},\mathbf{h}_{i}^{Q}\rangle$ $|i\in D_{\text{train}}\},$ (7)
    $\mathcal{I}^{-}_{NC}=\{\langle\mathbf{h}_{i}^{V},\mathbf{h}_{j}^{Q}\rangle$ $|i\neq j,i\in D_{\text{train}},j\in D_{\text{train}}\}.$ (8)
    And we define the positive pairs and negative pairs for CG alignment purpose as:
    $\mathcal{I}^{+}_{CG}=\{\langle\phi_{1}(c_{i}),\phi_{2}(g_{i})\rangle|$ $i\in D_{\text{train}}\},$ (9)
    $\mathcal{I}^{-}_{CG}=\{\langle\phi_{1}(c_{i}),\phi_{2}(g_{j})\rangle|$ $i\neq j,i\in D_{\text{train}},j\in D_{\text{train}}\}.$ (10)
  • Structure Preserving Contrastive Learning. To preserve the structural information of the graphical views, we perform intra-modality contrastive learning among the graphical views and their corrupted views. Concretely, we corrupt each of the graphical view $g_{i}$ with the edge dropping operation to obtain its corrupted view $g_{i}^{\prime}$. The positive pairs and negative pairs for structure-preserving purpose are then designed as:
    $\mathcal{I}^{+}_{\text{preserve}}$ =$=\{\langle\phi_{2}(g_{i}),\phi_{2}(g_{i}^{\prime})\rangle|i\in D_{\text{train}}\},$ (11)
    $\mathcal{I}^{-}_{\text{preserve}}$ =$=\{\langle\phi_{2}(g_{i}),\phi_{2}(g_{j}^{\prime})\rangle|i\neq j,i\in D_{\text{train}},j\in D_{\text{train}}\}.$ (12)
To help the backbone LLMs to digest the graphical views, we tune the LLMs with the expert soft signal. The prompt for finetuning consists of the system prompt, retrieved knowledge where the expert encoded graphical view is contained using a token embedding, and task prompt, as illustrated in Figure 4 (b). After the finetuning stage, we used the tuned models to generate codes using the soft prompting template.
@@ 3 Experiments
    Can soft prompting enhance the capability of the backbone LLMs? Does finetuning with the expert signals outperform the simple supervised finetuning ?
@@
    How is the compatibility of the graph extraction process?
3.1 Setup
In this paper, we evaluate CodeGRAG with the widely used HumanEval-X [^2] https://github.com/THUDM/CodeGeeX Zheng et al. (2023) dataset, competition-style CodeForce [^3] https://paperswithcode.com/sota/code-generation-on-apps and APPS [^4] https://arxiv.org/pdf/2105.09938v3 Hendrycks et al. (2021) datasets. We use greedy decoding strategy to do the generation. The evaluation metric is Pass@1. More details of the retrieval pool and the finetuning settings can be found in Section A.1 in the Appendix.